Given that on one section of the SAT the mean is 500 and the standard deviation is 100, what is the approximate probability of selecting a student whose score is 400 or lower followed by selecting a student whose score is 500 or higher on the test?
Answer #1

Solution :

Given that ,

a) P(x < 400)

= P[(x - ) / < (400 - 500) / 100]

= P(z < -1.00)

Using z table,

= 0.1587

= 0.16

b)P(x > 500) = 1 - p( x< 500 )

=1- p P[(x - ) / < (500 - 500) /100 ]

=1- P(z < 0)

= 1 - 0.50

= 0.50